Mr. P's hybrid car costs 6 cents a mile to operate, and cost $25,000 to buy. Ms. Sullivan's car costs $20,000 to buy and cost 10 cents a mile to operate. Write an expression or equation for the cost to operate each car as a function with respect to the number of miles driven. When will the cars have cost the same amount to operate?
Answer by checkley71(8403)

25000+.06X=20000+.10X
.06X-.10X=20000-25000
-.04X=-5000
X=-5000/-.04
X=125000 MILES DRIVEN
PROOF
25000+.06*125000=20000+.10*125000
25000+7500=20000+12500
32500=32500